OpenStudy (anonymous):

Sketch the solution to the initial value problem and determine its maximum value.

OpenStudy (anonymous):

\[dy/dt=2y-2yt\]

OpenStudy (anonymous):

y(0)=3

OpenStudy (experimentx):

\[ \frac{dy}{dt} = 2y(1-t)\] \[ \frac{dy}{y} = 2(1-t)dt\] Integrating both sides, \( \ln y = 2(t - t^2/2) +C\) , put the value 0,3 you get, c = ln(3) put the value of C ... and get your answer